Hi! Can you please help me with this problem.
The length of a rectangular road sign is 2 feet less than three times its width. Find the dimensions of the sign if the perimeter is 28 feet.
Thank you so much!

Answer by rapaljer(4671)

Let x= width
3x-2= length

2(x)+2(3x-2)=28
2x+6x-4=28
8x-4=28
8x=32
x=4 feet Width
3x-2=3*4-2=10 feet Length

Check: 2*4+2*10=8+20=28 It checks!!

For additional help with WORD PROBLEMS, please see my OWN website by clicking on my tutor name "rapaljer" anywhere in algebra.com. Click on my website which is given there, and look on my Homepage for the link "Basic, Intermediate, and College Algebra: One Step at a Time." Select "Basic Algebra", and look in "Chapter 1" for "1.10 Word Problems Part I and II." In addition to a detailed explanation, with examples, exercises, and ALL the answers, many of the exercises are solved IN COLOR on the "Math in Living Color" pages that go with these sections.

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida
Altamonte Springs Campus